# Course Allocation with Credits via Stable Matching

### v1

@@ 2 Model
    $c_{j}$ has fewer students assigned to it than its upper quota $q^{+}(c_{j})$ or $c_{j}$ prefers $s_{i}$ over some $s_{k}$ assigned to $c_{j}$. That is, $|S_{M}(c_{j})|<q^{+}(c_{j})$ or $s_{i}\succ_{c_{j}}s_{k}$ for some $s_{k}\in S_{M}(c_{j})$.
@@
    There exists a subset $D_{i}\subseteq C_{M}(s_{i})$ (possibly empty) such that $s_{i}$ prefers $c_{j}$ to every $c_{k}\in D_{i}$, $O(c_{j})\geq O(D_{i})$, and $O(C_{M}(s_{i}))-O(D_{i})+O(c_{j})\leq T(s_{i})$. That is, $s_{i}$ has enough credits remaining for $c_{j}$, possibly after removing a subset of courses that $s_{i}$ prefers less than $c_{j}$ and with a total number of credits smaller than or equal to that of $c_{j}$.
@@
    If there are downward-feasible constraints, the matching $(M\cup\{(s_{i},c_{j})\})\setminus\{(s_{i},c_{k}):c_{k}\in D_{i}\}$ is feasible. That is, removing $D_{i}$ and adding $c_{j}$ to $C_{M}(s_{i})$ does not result in an infeasible matching.
If $M$ has no size-blocking pair then it is pair-size-stable. Course Allocation with pair-size stability is denoted ca-ps.
The only difference between pair stability and pair-size stability is that we require that a student in a size-blocking pair must not end up with fewer credits than they started with when participating in a size-blocking pair. However, this may have unintended consequences. For example, consider agent $s_{1}$ from Figure 2. If $s_{1}$ is matched with $c_{3}$, then she would prefer to be matched with $c_{1}$ and $c_{2}$ instead. However, because $O(c_{3})=2$ and $O(c_{1})=O(c_{2})=1$, $s_{1}$ cannot form a size-blocking pair with either $c_{1}$ or $c_{2}$. This motivates our next stability definition, in which we allow students to block with more than one course.
@@
[…]
@@
Figure 2: Example of a pair-stable matching in a Course Allocation instance, highlighted in green. A subset of courses that $s_{1}$ may prefer over $c_{3}$ is shown in red.
@@
    for each $c_{j}\in B$, either $c_{j}$ has fewer students assigned to it than its upper quota $q^{+}(c_{j})$ or $c_{j}$ prefers $s_{i}$ over some $s_{k}$ assigned to $c_{j}$. That is, for every $c_{j}\in B$, either $|S_{M}(c_{j})|<q^{+}(c_{j})$ or $s_{i}\succ_{c_{j}}s_{k}$ for some $s_{k}\in S_{M}(c_{j})$.
@@
    There exists a subset $D_{i}\subseteq C_{M}(s_{i})$ (possibly empty) such that $s_{i}$ prefers every $c_{j}\in B$ to every $c_{k}\in D_{i}$, $O(B)\geq O(D_{i})$ and $O(C_{M}(s_{i}))-O(D_{i})+O(B)\leq T(s_{i})$. That is, $s_{i}$ has enough credits remaining for all courses in $B$, possibly after removing a subset of courses that $s_{i}$ prefers less than all courses in $B$ and with a total number of credits smaller than or equal to that of $B$.
@@
    If there are downward-feasible constraints, the matching $(M\cup\{(s_{i},c_{j})|c_{j}\in B\})\setminus\{(s_{i},c_{k}):c_{k}\in D_{i}\}$ is feasible. That is, removing $D_{i}$ and adding $c_{j}$ to $C_{M}(s_{i})$ for every $c_{j}\in B$ does not result in an infeasible matching.
If $M$ has no blocking coalition then it is coalition-stable. Course Allocation with coalition stability is denoted ca-c.
Under coalition stability, $s_{1}$ from Figure 2 can form a blocking coalition with $c_{1}$ and $c_{2}$ if she is matched with $c_{3}$. Moving on, we now consider at Figure 3. Here, if $s_{1}$ is matched with $c_{3}$, then she cannot form a blocking coalition with $c_{1}$ and $c_{2}$, because she prefers $c_{3}$ over $c_{2}$. However, it is plausible that some students would rather be matched with $c_{1}$ and $c_{2}$ instead, because their best course in the new matching is strictly better than the best course in the old matching. Hence, we obtain a new definition of stability.
@@
[…]
@@
Figure 3: Example of a coalition-stable matching in a Course Allocation instance, highlighted in green. A subset of courses that $s_{1}$ may prefer over $c_{3}$ is shown in red.
@@
    for each $c_{j}\in B$, either $c_{j}$ has fewer students assigned to it than its upper quota $q^{+}(c_{j})$ or $c_{j}$ prefers $s_{i}$ over some $s_{k}$ assigned to $c_{j}$. That is, for every $c_{j}\in B$, either $|S_{M}(c_{j})|<q^{+}(c_{j})$ or $s_{i}\succ_{c_{j}}s_{k}$ for some $s_{k}\in S_{M}(c_{j})$.
@@
    There exists a subset $D_{i}\subseteq C_{M}(s_{i})$ (possibly empty) such that $s_{i}$ prefers her most preferred course in $B$ to her most preferred course in $D_{i}$, $O(B)\geq O(D_{i})$ and $O(C_{M}(s_{i}))-O(D_{i})+O(B)\leq T(s_{i})$. That is, $s_{i}$ has enough credits remaining for all courses in $B$, possibly after removing a subset of courses $D_{i}$ with a total number of credits smaller than or equal to that of $B$ such that $s_{i}$ prefers her most preferred course in $B$ over her most preferred course in $D_{i}$.
@@
    If there are downward-feasible constraints, the matching $(M\cup\{(s_{i},c_{j})|c_{j}\in B\})\setminus\{(s_{i},c_{k}):c_{k}\in D_{i}\}$ is feasible. That is, removing $D_{i}$ and adding $c_{j}$ to $C_{M}(s_{i})$ for every $c_{j}\in B$ does not result in an infeasible matching.
If $M$ has no first-blocking coalition then it is first-coalition-stable. Course Allocation with first-coalition stability is denoted ca-fc.
@@
We will prove all three implications by contradiction. First, assume that a matching $M$ is pair-stable but not first-coalition-stable. Hence, there exist a student $s_{i}$ and a set of courses $B$ such that $s_{i}$ has enough space for all courses in $B$, possibly after removing a subset of courses $D_{i}\subseteq C_{M}(s_{i})$ that $s_{i}$ prefers less than her most preferred course in $B$. Let this course be $c_{j}$. Then $(s_{i},c_{j})$ is a blocking pair, so $M$ is not pair-stable, a contradiction. If there are downward-feasible constraints, $(s_{i},c_{j})$ is still a blocking pair, because $s_{i}$ and $B$ form a first-blocking coalition, which means that $(M\cup\{(s_{i},c_{j})|c_{j}\in B\})\setminus\{(s_{i},c_{k}):c_{k}\in D_{i}\}$ is feasible, and $(M\cup(s_{i},c_{j}))\setminus\{(s_{i},c_{k}):c_{k}\in D_{i}\}$ is a subset of that matching, so also feasible.
Now, assume that $M$ is first-coalition-stable, but not coalition-stable. Then there exist a student $s_{i}$ and a set of courses $B$ such that $s_{i}$ has enough space for all courses in $B$, possibly after removing a subset of courses $D_{i}\subseteq C_{M}(s_{i})$ that $s_{i}$ prefers less than all courses in $B$. This implies that $s_{i}$ prefers her most preferred course in $B$ over her most preferred course in $D_{i}$, so $s_{i}$ and $B$ form a first-blocking coalition, and $M$ is not first-coalition-stable, a contradiction. If there are downward-feasible constraints, $s_{i}$ and $B$ is still a first-blocking coalition, because $s_{i}$ and $B$ form a blocking coalition, which means that $(M\cup\{(s_{i},c_{j})|c_{j}\in B\})\setminus\{(s_{i},c_{k}):c_{k}\in D_{i}\}$ is feasible.
Finally. assume that $M$ is coalition-stable, but not pair-size-stable. Then there exist a student $s_{i}$ and a course $c_{j}$ such that $(s_{i},c_{j})$ is a size-blocking pair. However, this implies that $s_{i}$ and $\{c_{j}\}$ form a blocking coalition, so $M$ is not coalition-stable, a contradiction. If there are downward-feasible constraints, $s_{i}$ and $\{c_{j}\}$ is still a blocking coalition, because $(s_{i},c_{j})$ is a size-blocking pair, which means that $(M\cup(s_{i},c_{j}))\setminus\{(s_{i},c_{k}):c_{k}\in D_{i}\}$ is feasible. ∎
@@ 3 Verifying stability
In this section we consider the complexity of the problem of determining whether a given matching is stable. Firstly, we remark that it is fairly trivial to check whether an assignment is a matching, and whether a matching is feasible, in the presence of downward-feasible constraints. One simply needs to check if the assignment respects upper quotas and credit limits, and if it respects all downward-feasible constraints. On the other hand, checking whether a (feasible) matching is stable may be trickier, depending on the stability definition. Let ca-x-test be the problem of deciding whether a given matching $M$ is stable, for a given definition of stability x. Under pair stability, one can verify the stability of $M$ in polynomial time. For any acceptable student-course pair $(s_{i},c_{j})\notin M$, we can check if $c_{j}$ has enough capacity for $s_{i}$, or if $c_{j}$ prefers $s_{i}$ over some $s_{k}\in S_{M}(c_{j})$, which takes $O(n)$ time, where $n$ is the number of agents. Similarly, we can check in $O(m)$, where $m$ is the number of courses, if $s_{i}$ has enough remaining credits for $c_{j}$, possibly after removing all courses in $C_{M}(s_{i})$ that $s_{i}$ prefers less than $c_{j}$. Recall that under this definition a student can end up with fewer credits after satisfying a blocking pair. As there are $O(nm)$ acceptable pairs, the total complexity is $O(n^{2}m^{2})$. Thus, we obtain the following observation.
@@
ca-p-test is solvable in $O(n^{2}m^{2})$ time, where $n=|S|$ and $m=|C|$.
However, verifying stability is hard for the other three definitions of stability, which we will show in Theorem 1. The difference is that in a blocking pair (in pair stability), a student is allowed to end up with fewer credits than she had in the matching $M$, but that is not allowed in the other three definitions of stability. This fact is crucial in our transformation from the Subset Sum problem. An instance of Subset Sum consists of a multiset of integers $X$ and a target integer $T$. The goal is to determine the existence of a subset of $X$ where the sum of its elements equals $T$. The problem is weakly NP-complete, even if every element in $X$ is a positive integer (garey_1979_computers, , page 223).
@@
These problems are in co-NP because given a matching $M$, a student $s$, a set of courses $B$ (of size $1$ in the case of pair-size stability) with which $s$ forms a size-blocking pair/blocking coalition/first-blocking coalition, and a set of courses $D$ that $s$ would drop, we can check whether $s$ blocks the matching $M$ with the courses in $B$.
Consider a Subset Sum instance $I$, with multiset $X$ and target number $T$. We create an instance $J$ of Course Allocation as follows. For every element $e\in X$, we create a course $c_{e}$, with upper quota $1$ and $O(c_{e})=e$. We create one additional course, $b$, with upper quota $1$ and $O(b)=T$. We also create one student, $s$, with $T(s)$ equal to the sum of all elements in $X$. The preference list of $s$ is $b$ first, then all of the other courses in any order. The matching that we want to verify is $M=\{(s,c_{e})|e\in X\}$. Thus, the matching includes every acceptable pair except for $(s,b)$. Also, $O(C_{M}(s)=T(s))$.
Assume that there exists a subset $X^{\prime}\subseteq X$ such that $\sum_{e\in X^{\prime}}e=T$. Let $D=\{c_{e}|e\in X^{\prime}\}$ be a subset of courses. Then $(s,b)$ is a size-blocking pair, because $b$ has enough capacity for $s$, $s$ prefers $b$ over every course in $D$, $O(b)\geq O(D)$, and $O(C_{M}(s))-O(D)+O(b)\leq T(s)$. Thus, $M$ is not pair-size-stable. We have that $s$ and $\{B\}$ form a blocking coalition and a first-blocking coalition, too, so $M$ is not coalition-stable or first-coalition-stable either.
Now, assume that $M$ is not pair-size-stable. Thus, there exists at least one size-blocking pair. The only acceptable pair that is not in $M$ is $(s,b)$, so that is our size-blocking pair. As $(s,b)$ is a size-blocking pair, $s$ prefers $b$ over a subset of courses $D\subseteq C_{M}(s)$ such that $O(b)\geq O(D)$ and $O(C_{M}(s))-O(D)+O(b)\leq T(s)$. As $T(s)-O(C_{M}(s))+O(D)=O(b)$, we have that $O(b)=T=O(D)$. As $b\notin D$, all courses in $D$ are of type $c_{e}$. Thus, we can build a subset $X^{\prime}\subseteq X$, such that $X=\{e|c_{e}\in D\}$ and $\sum_{e\in X^{\prime}}e=T$. Therefore, there exists a subset of $X$ whose elements add up to $T$. This direction also works in the case of coalition stability or first-coalition stability. ∎
However, just like with Subset Sum, a pseudo-polynomial dynamic programming algorithm can verify the stability of a given matching, for either pair-size stability, coalition stability, or first-coalition stability. Thus, if the number of credits of all courses is not exponentially large, verifying the stability of a matching can be done in polynomial time. In real-life settings, it is reasonable to assume that course credits are polynomially bounded.

### v2

@@ 2 Model
    Either $c_{j}$ is undersubscribed or $c_{j}$ prefers $s_{i}$ over some $s_{k}$ assigned to $c_{j}$. That is, $|S_{M}(c_{j})|<q^{+}(c_{j})$ or $s_{i}\succ_{c_{j}}s_{k}$ for some $s_{k}\in S_{M}(c_{j})$.
@@
    There exists a subset $D_{i}\subseteq C_{M}(s_{i})$ (possibly empty) such that $s_{i}$ prefers $c_{j}$ to every $c_{k}\in D_{i}$, $O(c_{j})\geq O(D_{i})$, and $O(C_{M}(s_{i}))-O(D_{i})+O(c_{j})\leq T(s_{i})$. That is, $s_{i}$ has enough credits remaining for $c_{j}$, possibly after removing a subset of courses that $s_{i}$ prefers $c_{j}$ to and with a total number of credits less than or equal to that of $c_{j}$.
@@
    If there are downward-feasible constraints, $(C_{M}(s_{i})\setminus D_{i})\cup\{c_{j}\}\in F_{i}$. That is, removing $D_{i}$ and adding $c_{j}$ to $C_{M}(s_{i})$ does not result in an infeasible matching.
If $M$ has no size-blocking pair then it is pair-size-stable. ca with pair-size stability is denoted ca-ps.
The only difference between pair stability and pair-size stability is that we require that a student in a size-blocking pair must not end up with fewer credits than they started with when participating in a size-blocking pair. However, this may have unintended consequences. For example, consider student $s_{1}$ from Figure 2. If $s_{1}$ is matched with $c_{3}$, then she would prefer to be matched with $c_{1}$ and $c_{2}$ instead. However, because $O(c_{3})=2$ and $O(c_{1})=O(c_{2})=1$, $s_{1}$ cannot form a size-blocking pair with either $c_{1}$ or $c_{2}$. This motivates our next stability definition, in which we allow students to block with more than one course.
@@
$s_{1}$ 2 $c_{1}$ $\succ$ $c_{2}$ $\succ$ $\boxed{c_{3}}$ $c_{1}$ 1 1 $s_{2}$ $\succ$ $s_{1}$
$s_{2}$ 1 $\boxed{c_{2}}$ $\succ$ $c_{1}$ $c_{2}$ 1 1 $s_{1}$ $\succ$ $\boxed{s_{2}}$
@@
Figure 2: Example of a pair-size-stable matching in a ca instance, highlighted in green. A subset of courses that $s_{1}$ may prefer over $c_{3}$ is shown in red.
@@
    for each $c_{j}\in B$, either $c_{j}$ is undersubscribed or $c_{j}$ prefers $s_{i}$ over some $s_{k}$ assigned to $c_{j}$. That is, for every $c_{j}\in B$, either $|S_{M}(c_{j})|<q^{+}(c_{j})$ or $s_{i}\succ_{c_{j}}s_{k}$ for some $s_{k}\in S_{M}(c_{j})$.
@@
    There exists a subset $D_{i}\subseteq C_{M}(s_{i})$ (possibly empty) such that $s_{i}$ prefers every $c_{j}\in B$ to every $c_{k}\in D_{i}$, $O(B)\geq O(D_{i})$ and $O(C_{M}(s_{i}))-O(D_{i})+O(B)\leq T(s_{i})$. That is, $s_{i}$ has enough credits remaining for all courses in $B$, possibly after removing a subset of courses that $s_{i}$ prefers all courses in $B$ to, and with a total number of credits less than or equal to that of $B$.
@@
    If there are downward-feasible constraints, $(C_{M}(s_{i})\setminus D_{i})\cup B\in F_{i}$. That is, removing $D_{i}$ and adding $B$ to $C_{M}(s_{i})$ does not result in an infeasible matching.
If $M$ has no blocking coalition then it is coalition-stable. ca with coalition stability is denoted ca-c.
Under coalition stability, $s_{1}$ from Figure 2 can form a blocking coalition with $c_{1}$ and $c_{2}$ if she is matched with $c_{3}$. Moving on, we now consider Figure 3. Here, if $s_{1}$ is matched with $c_{3}$, then she cannot form a blocking coalition with $c_{1}$ and $c_{2}$, because she prefers $c_{3}$ over $c_{2}$. However, it is plausible that $s_{1}$ would rather be matched with $c_{1}$ and $c_{2}$ instead, because her best course in the new matching is strictly better than her best course in the old matching. Hence, we obtain a new definition of stability.
@@
$s_{1}$ 2 $c_{1}$ $\succ$ $\boxed{c_{3}}$ $\succ$ $c_{2}$ $c_{1}$ 1 1 $s_{2}$ $\succ$ $s_{1}$
$s_{2}$ 1 $\boxed{c_{2}}$ $\succ$ $c_{1}$ $c_{2}$ 1 1 $s_{1}$ $\succ$ $\boxed{s_{2}}$
@@
Figure 3: Example of a coalition-stable matching in a ca instance, highlighted in green. A subset of courses that $s_{1}$ may prefer over $c_{3}$ is shown in red.
@@
    for each $c_{j}\in B$, either $c_{j}$ is undersubscribed or $c_{j}$ prefers $s_{i}$ over some $s_{k}$ assigned to $c_{j}$. That is, for every $c_{j}\in B$, either $|S_{M}(c_{j})|<q^{+}(c_{j})$ or $s_{i}\succ_{c_{j}}s_{k}$ for some $s_{k}\in S_{M}(c_{j})$.
@@
    There exists a subset $D_{i}\subseteq C_{M}(s_{i})$ (possibly empty) such that $s_{i}$ prefers her most-preferred course in $B$ to her most-preferred course in $D_{i}$, $O(B)\geq O(D_{i})$ and $O(C_{M}(s_{i}))-O(D_{i})+O(B)\leq T(s_{i})$. That is, $s_{i}$ has enough credits remaining for all courses in $B$, possibly after removing a subset of courses $D_{i}$ with a total number of credits less than or equal to that of $B$ such that $s_{i}$ prefers her most-preferred course in $B$ over her most-preferred course in $D_{i}$.
@@
    If there are downward-feasible constraints, $(C_{M}(s_{i})\setminus D_{i})\cup B\in F_{i}$. That is, removing $D_{i}$ and adding $B$ to $C_{M}(s_{i})$ does not result in an infeasible matching.
If $M$ has no first-blocking coalition then it is first-coalition-stable. ca with first-coalition stability is denoted ca-fc.
@@
We will prove all three implications by contradiction. First, assume that a matching $M$ is pair-stable but not first-coalition-stable. Hence, there exist a student $s_{i}$ and a set of courses $B$ such that $s_{i}$ has enough space for all courses in $B$, possibly after removing a subset of courses $D_{i}\subseteq C_{M}(s_{i})$ that $s_{i}$ prefers her most-preferred course in $B$ to. Let this course be $c_{j}$. Then $(s_{i},c_{j})$ is a blocking pair, so $M$ is not pair-stable, a contradiction. If there are downward-feasible constraints, $(s_{i},c_{j})$ is still a blocking pair, because $s_{i}$ and $B$ form a first-blocking coalition, which means that $(C_{M}(s_{i})\setminus D_{i})\cup B\in F_{i}$. Moreover $(C_{M}(s_{i})\setminus D_{i})\cup\{c_{j}\}\subseteq(C_{M}(s_{i})\setminus D_{i})\cup B$ and hence $(C_{M}(s_{i})\setminus D_{i})\cup\{c_{j}\}\in F_{i}$.
Now, assume that $M$ is first-coalition-stable, but not coalition-stable. Then there exist a student $s_{i}$ and a set of courses $B$ such that $s_{i}$ has enough space for all courses in $B$, possibly after removing a subset of courses $D_{i}\subseteq C_{M}(s_{i})$ such that $s_{i}$ prefers each course in $B$ to each course in $D_{i}$. This implies that $s_{i}$ prefers her most-preferred course in $B$ over her most-preferred course in $D_{i}$, so $(s_{i},B$) is a first-blocking coalition, and $M$ is not first-coalition-stable, a contradiction. If there are downward-feasible constraints, $(s_{i},B$) is still a first-blocking coalition, since $(s_{i},B$) is a blocking coalition, which means that $(C_{M}(s_{i})\setminus D_{i})\cup B\in F_{i}$.
Finally, assume that $M$ is coalition-stable, but not pair-size-stable. Then there exist a student $s_{i}$ and a course $c_{j}$ such that $(s_{i},c_{j})$ is a size-blocking pair. However, this implies that $(s_{i},\{c_{j}\})$ is a blocking coalition, so $M$ is not coalition-stable, a contradiction. If there are downward-feasible constraints, $(s_{i},\{c_{j}\})$ is still a blocking coalition, because $(s_{i},c_{j})$ is a size-blocking pair, which means that $(C_{M}(s_{i})\setminus D_{i})\cup\{c_{j}\}\in F_{i}$. ∎
@@ 3 Verifying stability
In this section we consider the complexity of the problem of determining whether a given matching is stable, with respect to the stability criteria defined in the previous section. Firstly, we remark that it is straightforward to check whether an assignment is a matching, and whether a matching is feasible, in the presence of downward-feasible constraints. One simply needs to check if the assignment respects upper quotas and credit limits, and if it respects all downward-feasible constraints. On the other hand, checking whether a (feasible) matching is stable may be trickier, depending on the stability definition. Let ca-x-test be the problem of deciding whether a given matching $M$ is stable, for a given definition of stability x. Under pair stability, one can verify the stability of $M$ in polynomial time. For each course we can compute its worst assigned student in $M$, within an overall time complexity of $O(L)$, where $L$ is the total number of acceptable pairs. For any acceptable student-course pair $(s_{i},c_{j})\notin M$, we can check if $c_{j}$ has enough capacity for $s_{i}$, or if $c_{j}$ prefers $s_{i}$ over its worst assigned student in $M$, which can be done in $O(1)$ time, using suitable data structures. Similarly, we can check in $O(m)$ time, where $m$ is the number of courses, if $s_{i}$ has enough remaining credits for $c_{j}$, possibly after removing all courses in $C_{M}(s_{i})$ that $s_{i}$ prefers $c_{j}$ to. Recall that under this definition a student can end up with fewer credits after satisfying a blocking pair. The overall time complexity is then $O(Lm)$. Thus, we obtain the following observation.
@@
ca-p-test is solvable in $O(Lm)$ time, where $m=|C|$ and $L$ is the number of acceptable pairs.
However, verifying stability is hard for the other three definitions of stability, which we will show in Theorem 1. The difference is that in a blocking pair (for pair stability), a student is allowed to end up with fewer credits than she had in the matching $M$, but that is not allowed in the other three definitions of stability. This fact is crucial in our transformation from the subset sum problem. An instance of subset sum consists of a multiset of elements $X$, where each element $e\in X$ has size $s(e)$, and a target integer $T$. The goal is to determine the existence of a subset of $X$ where the sum of the sizes of its elements equals $T$. The problem is weakly NP-complete, even if the size of every element in $X$ is a positive integer (Garey and Johnson, 1979, page 223).
@@
These problems are in co-NP because given a matching $M$, a student $s$, a set of courses $B$ (of size $1$ in the case of pair-size stability) with which $s$ would form a size-blocking pair/blocking coalition/first-blocking coalition, and a set of courses $D$ that $s$ would drop, we can check whether $s$ blocks the matching $M$ with the courses in $B$ in polynomial time.
Consider a subset sum instance $I$, with multiset $X$ and target number $T$. We create an instance $J$ of ca as follows. For every element $e\in X$, we create a course $c_{e}$, with upper quota $1$ and $O(c_{e})=s(e)$. We create one additional course, $b$, with upper quota $1$ and $O(b)=T$. We also create one student, $s$, with $T(s)$ equal to the sum of the sizes of all elements in $X$. The preference list of $s$ is $b$ first, then all of the other courses in any order. The matching that we want to verify is $M=\{(s,c_{e})|e\in X\}$. Thus, the matching includes every acceptable pair except for $(s,b)$. Also, $O(C_{M}(s))=T(s)$.
Assume that there exists a subset $X^{\prime}\subseteq X$ such that $\sum_{e\in X^{\prime}}s(e)=T$. Let $D=\{c_{e}|e\in X^{\prime}\}$ be a subset of courses. Then $(s,b)$ is a size-blocking pair, because $b$ has enough capacity for $s$, $s$ prefers $b$ over every course in $D$, $O(b)=O(D)=T$, and $O(C_{M}(s))-O(D)+O(b)=T(s)$. Thus, $M$ is not pair-size-stable. We have that $(s,B)$ is a blocking coalition and a first-blocking coalition, too, so $M$ is not coalition-stable or first-coalition-stable either.
Now, assume that $M$ is not pair-size-stable. Thus, there exists at least one size-blocking pair. The only acceptable pair that is not in $M$ is $(s,b)$, so that is our size-blocking pair. As $(s,b)$ is a size-blocking pair, $s$ prefers $b$ over a subset of courses $D\subseteq C_{M}(s)$ such that $O(b)\geq O(D)$ and $O(C_{M}(s))-O(D)+O(b)\leq T(s)$. As $O(C_{M}(s))=T(s)$, we have that $O(b)\leq O(D)$, which combined with $O(b)\geq O(D)$ gives us $O(b)=T=O(D)$. As $b\notin D$, all courses in $D$ are of type $c_{e}$. Thus, we can build a subset $X^{\prime}\subseteq X$, such that $X=\{e|c_{e}\in D\}$ and $\sum_{e\in X^{\prime}}s(e)=T$. Therefore, there exists a subset of $X$ whose elements’ sizes add up to $T$. This direction also works in the case of coalition stability or first-coalition stability. ∎
However, just as in the case of subset sum, a pseudo-polynomial dynamic programming algorithm can verify the stability of a given matching, for either pair-size stability, coalition stability, or first-coalition stability. Thus, if the number of credits of all courses is not exponentially large, verifying the stability of a matching can be done in polynomial time. In real-life settings, it is reasonable to assume that course credits are polynomially bounded. [^1] For example, all courses taken by Honours students at the School of Law, University of Glasgow Graves and Wierzbicki (2025), carry 20 or 40 credits. For the purposes of a dynamic programming algorithm, each course’s number of credits, and each student’s maximum number of credits, can be divided by 20.